Find the volume of the solid formed by revolving the region bounded by y=(x-2)² and y=x about the y -axis.


1
Expert's answer
2021-05-31T06:39:27-0400

Find the points of intersection of the parabola with the y=x

"x = (x-2)^2"

We get "x = 1,4"

As the region is revolved about the y axis, we express the equation of the bounding curve in terms of y


"y = (x-2)^2"


"(x-2) =\pm \sqrt{y}"


"x = 2 \pm \sqrt{y}"


Volume can be calculated as:


"V = \pi \int_{1}^{4}[(2+\sqrt{y})^2-(2-\sqrt{y})^2]dy"


"V = \pi \int_{1}^{4}4\sqrt{y}dy"


"V = \dfrac{8\pi}{3}[y^{3/2}]_{1}^{4}"


"V = \dfrac{56\pi}{3}"